A point charge of q = 4µC is 3cm apart from the charge q 1 = 1µC.

a. Find the magnitude of the Coulomb force which one particle exerts on the other.

b. Is the force attractive or repulsive?


1
Expert's answer
2022-01-21T08:25:16-0500

a) The magnitude of the Coulomb force which one particle exerts on the other is given as follows:


"F = k\dfrac{q_1q_2}{r^2}\\
F = 9\times 10^9\cdot \dfrac{4\times 10^{-6}\cdot 1\times 10^{-6}}{(0.03)^2} =40N"

b) The force is repulsive, since the particles have the charge of the same kind.